OpenStudy (sleepyjess):

@Whitemonsterbunny17

whitemonsterbunny17 (whitemonsterbunny17):

Imma hereee

OpenStudy (sleepyjess):

lol http://prntscr.com/52iyqc

whitemonsterbunny17 (whitemonsterbunny17):

My name is eurywhere!! Aaaaaahhhhh!!! DX Haha

OpenStudy (sleepyjess):

ok so first question In pentagon PQRST, the measure of angle P=60, and the measure of angle Q=130. Angle S and angle T are each three times as large as angle R. a. Find the measures of angle R, angle S, and angle T. b. Which pair of sides of PQRST must be parallel?

OpenStudy (sleepyjess):

6-+130+R+S+T=540

OpenStudy (sleepyjess):

whoops should be 60 not 6-

OpenStudy (sleepyjess):

-190 from each side gives us R+S+T=350

OpenStudy (sleepyjess):

S=3R T=3R so we know that S+T+R=7R

whitemonsterbunny17 (whitemonsterbunny17):

o.O

OpenStudy (sleepyjess):

following me so far?

whitemonsterbunny17 (whitemonsterbunny17):

So a is 7R?

OpenStudy (sleepyjess):

almost 7R=350 /7

whitemonsterbunny17 (whitemonsterbunny17):

50?

OpenStudy (sleepyjess):

yes R=50 S=3R T=3R so what is the measure of S and T?

whitemonsterbunny17 (whitemonsterbunny17):

S=350/3, right?

OpenStudy (sleepyjess):

no just 3*50

whitemonsterbunny17 (whitemonsterbunny17):

Ohh xD 150?

OpenStudy (sleepyjess):

yes

OpenStudy (sleepyjess):

not exactly sure how you would make the shape though

whitemonsterbunny17 (whitemonsterbunny17):

Hrm, so S is parallel to T, right?

OpenStudy (sleepyjess):

O.o http://prntscr.com/52j11t

OpenStudy (sleepyjess):

2 different people

whitemonsterbunny17 (whitemonsterbunny17):

Yea, people do that xD

OpenStudy (wolf1728):

Whitemonsterbunny17 I came here about the previous problem you posted about the quadrilateral.

whitemonsterbunny17 (whitemonsterbunny17):

Ohh, yes, what about it? ^.^

OpenStudy (wolf1728):

I finished my diagram and posted it - I think thank no sides are parallel

OpenStudy (wolf1728):

think THAT

whitemonsterbunny17 (whitemonsterbunny17):

I saw it and I did reply, but why does the book ask directly for what the parallel sides are if none exist?

OpenStudy (wolf1728):

I don't know. Sometimes books, teachers, etc like to ask questions like that. Things like what is the fourth side of a triangle called? (That sort of thing).

whitemonsterbunny17 (whitemonsterbunny17):

Ohh, okay. That makes sense I suppose. Thanks again! ^.^

OpenStudy (wolf1728):

ok you are welcome :-)

whitemonsterbunny17 (whitemonsterbunny17):

Okay, so back to this problem.. S is parallel to T, correct?

OpenStudy (sleepyjess):

ive tried multiple times and i just cant get this shape correct @wolf1728

OpenStudy (sleepyjess):

with as close as i can get it, it would appear that PQ and RS are parrellel